# Automation of Quantum Dot Measurement Analysis via Explainable Machine Learning

### v4

@@ 2.5 Synthetic triangle fitting approach
In practice, this involves applying an ideal low-pass filter, such as a Gaussian filter. To focus on the details of the structure, we use the magnitude gradient of the low-pass filtered image. While this can be thought of as a simple edge detection, we do not have a priori knowledge about the appropriate scale of the Gaussian for any particular image. To quantify the similarity between an experimentally acquired and synthetic triangle plot, we use the $L_{2}$-norm applied to both thus transformed images, assuming that two images are visually similar if their distance in this transformed space is small.
Definition 3
@@
$q^{*},\sigma^{*}=\arg\!\min_{q,\sigma}\big{\{}\lambda\,\mathcal{S}_{\sigma}^{DG}((\mathcal{I}_{\rm exp},\mathcal{I}_{\rm sim})+(1-\lambda)\,\mathcal{S}^{I}((\mathcal{I}_{\rm exp},\mathcal{I}_{\rm sim})+\epsilon(q)\big{\}},$ (6)
where $DG_{\sigma}(U)=\nabla(G_{\sigma}(U))$ is the gradient of the Gaussian transform; $I(U)=U$ is the identity transform; $\lambda$ is a hyperparameter giving us control over the balance between trusting $DG_{\sigma}$ and $I$ similarity measures; and
@@
for some arbitrarily large penalty constant $C$.
In an experimental context, we have access to additional data tied to each image to aid in establishing initial guesses for optimization. However, in this context we only rely solely on the raw image. To remedy this, we use the differential evolution global optimizer implemented in SciPy [34] This makes our method less sensitive to initial values than if we had used a local optimizer.
@@ 3 Results
We present the results of experiments with the Gabor filterbank approach, the synthetic triangle fitting approach, and a combination of the synthetic features and the most informative Gabor feature. The transformation procedures described in sections 2.3 and 2.5 are applied to each experimentally observed triangle to produce the final vector of features $\textbf{v}(\mathcal{I}_{exp})$. In each case, the vectorized features $\textbf{v}(\mathcal{I}_{exp})$ are used to train and test an EBM model.
To improve the accuracy and interpretability at a modest training time cost, we utilize the smoothing and greedy rounds parameters [35]. For a more thorough discussion of this, see Ref. [36]. We carry out five six-fold stratified cross-validations and report averaged results and type I and type II errors for each method. For completeness, we also include the averaged confusion matrices.
The results from all experiments, presented in table LABEL:tab:model_comparison, show a relatively close performance for all three methods, with the Gabor fitting approach only slightly overperforming the more intuitive synthetic triangle fitting approach, at $92.5(1.2)~{}\%$ vs. $90.5(1.0)~{}\%$, respectively. Interestingly, the inclusion of a single, most important Gabor filter in the vector of features obtained using synthetic triangle fitting brings the EMBs performance back up to $91.7(8)~{}\%$. This validates that the optimal fit captures the location and average intensity of the walls and diagonal components. Modifying the diagonal component function to capture the ridged pattern observed in the experimental data could further improve the performance.
Figure 3: The feature importance plot for (a) the Gabor filterbank vectorization approach, (b) the synthetic triangle fitting vectorization approach, and (c) the combined vectorization approach. The mean absolute score refers to the extent to which a feature contributes to all predictions. Feature curve for (d) the oriented Gabor feature $\mathcal{G}_{16,16}^{45}$, (e) the fit fitness feature $\mathcal{F}$ of the synthetic triangle fit, and (f) the diagonal boundary component $\mathcal{D}_{B}$ of the synthetic triangle fit for the synthetic triangle fitting approach shown in orange and for the mixed approach shown in blue. The $x$-axis indicates the value attained by the feature [normalized to $(0,1)$ in (d) and (e)], and the score indicates the log odds towards good (positive) and bad (negative) classes for each respective feature. The histograms in (g-i) depict the relative class distribution as a function of the normalized feature value, with cyan representing good triangles and pink representing bad triangles.
@@ 3.1 Prediction interpretability
The feature importance plot for the synthetic triangle fitting approach, shown in figure 3(b), reveals that this model also prioritizes features characterizing the triangle region, with $\mathcal{D}_{B}$ (the diagonal boundary component) and $\mathcal{D}_{M}$ (the diagonal magnitude component) being among the top three filters. The most important feature here is the fit fitness $\mathcal{F}$, quantifying how close the vectorized representation matches the experimental data. Looking at the feature curve for $\mathcal{F}$ depicted in figure 3(e), we see that for a modest cost associated with fitting, that is when $10^{2}\lessapprox\mathcal{F}\lessapprox 10^{3}$, the log odds of belonging to the good class are positive. For bad triangles, the associated cost is either disproportionately low or high. On the low end of the spectrum, this is likely due to the problem of fitting a bad triangle with only two rectangular regions being a significantly simpler optimization problem. On the high end, the exceptionally high cost is generally associated with poor quality of the data.
@@
The second most important feature in the synthetic triangle fitting approach is $\mathcal{D}_{B}$. The $\mathcal{D}_{B}$ feature curve, shown in orange in figure 4(f), indicates that the log odds that a triangle region belongs to a good plot increase with $\mathcal{D}_{B}$ once it extends past 28. The log odds of a triangle region belonging to a good plot are positive when $8535<\mathcal{D}_{B}<$ 85. The $\mathcal{D}_{B}$ feature quantifies the extent to which the triangle region stretches in the screening gate-screening gate plot, as depicted in figure 2(c). When $\mathcal{D}_{B}<35$, it is likely that the fitted triangle region overlaps with the horizontal and/or vertical sigmoid, and thus $\mathcal{D}_{B}$ would not be a reliable discriminant between good and bad class. While theoretically vector $\mathcal{D}_{B}$ can extend to the bottom left corner in figure 2(c), for large values of $\mathcal{D}_{B}$ the hypotenuse becomes relatively short which makes finding a reliable fit challenging. This dependency is confirmed by the feature plot, indicating that for $\mathcal{D}_{B}>85$, the log odds of a triangle region belonging to a good class fall below 00. The importance of the $\mathcal{D}_{B}$ filter is further evidenced by its dominant role in the combined approach, where it occupies the third most important position, see figure 4(c).
@@
It is important to emphasize that, due to the way EBMs are trained, the feature curves are not the same as the marginal distribution of feature values. Rather, each curve is affected by the complete vector of features $\textbf{v}(\mathcal{I}_{exp})$ used to train the model. This means that individual features cannot be treated as entirely independent from one another. This is confirmed by the slight difference between the shape of the $\mathcal{D}_{B}$ curve for the synthetic triangle fitting approach vs. the mixed approach, shown in figure 3(f) in orange and blue, respectively.
@@
The combined model achieves comparable performance to the Gabor-based model while including only a single Gabor feature in addition to the synthetic features. It thus retains the improved intelligibility of the synthetic features without a significant reduction in accuracy.
@@
Figure 4: (a) A sample experimentally acquired bad scan. (e, i) Two sample experimentally acquired good scans. (b, f, j) The fitted synthetic triangle plots for scans (a), (e), and (i), respectively. (c, g, k) Gabor-filter-based local interpretation of scans shown in (a), (e), and (i), respectively. (d, h, l) The feature importance for the combined vectorization method for scans (a), (e), and (i), respectively. The intercept coefficient $\alpha=4$ is removed from plots (d), (h), and (l) for clarity.
@@
EBMs also give us the ability to examine predictions for individual images. Figure 4(a), (b), and (c) shows an example of a bad image, its synthetic fit, and Gabor filter representation, respectively. The synthetic triangle fit effectively captures the structural content of the experimental image, exhibiting no triangle region. The relative feature importance plot, shown in figure 4(d), also agrees with this observation, indicating $\mathcal{V}_{B}$ (the vertical boundary component) and $\mathcal{H}_{B}$ (the horizontal boundary component) as the most dominant features contributing to the negative (i.e., bad class) prediction.
@@
In all three examples, there is a direct observable relationship between the fit parameters and the visual features of the synthetic triangle. In practical applications, these visual features can lead to thresholding decisions and derived quantities for choosing operating points in the voltage space. For example, in the case of figure 4(i), a good operating point might be right inside the triangle region above the diagonal boundary, bisecting the area of the triangle at around $(V_{S_{3}},V_{S_{2}})=(0.09~{}$\mathrm{V}$,0.09~{}$\mathrm{V}$)$ [at point (25,25) in figure 4(j)]. Using the $\mathcal{D}_{B}$ and $\mathcal{D}_{\theta}$ with the knowledge of the boundaries of the image in voltage space, along with the $\mathcal{V}_{B}$ and $\mathcal{H}_{B}$, we can create a normal vector describing the direction in which the triangle is expanding. This normal vector describes the relative strengths of the associated screening gates. Ideally, the normal vector of the triangle would point symmetrically towards the bottom left corner, meaning that both screening gates have equal action on the forming current channel. But, more often than not, this vector tends to vary. By re-scaling the normal vector and defining the magnitude to begin at ($\mathcal{V}_{B}$,$\mathcal{H}_{B}$) with the end pointing to the diagonal boundary of the triangle, the vector can now encode size information. If we then reduce the magnitude by $10~{}\%$, we have an initial guess, inside the bounds of the triangle, for the operating point. We can also use the vector to calculate the area of the triangle in voltage space which can be thresholded to make sure it is large enough to proceed.
@@ 4 Conclusion and outlook
In this work, we demonstrate an alternative approach to vectorizing image data for use with EBMs that relies on generating synthetic data that best fits the experimentally acquired scans. We show that the new method is better adapted to the complexity of the triangle plots dataset. While the original vectorization method, invoking the Gabor filterbank features, produces a model that performs well with the classification task, its interpretability and usefulness are limited as the resulting model does not comprehend the underlying structure of the data. The alternative method, relying on fitting synthetic triangle plots to the experimental data, retains comparable accuracy, failing in about $2~{}\%$ more cases than the Gabor filterbank approach. However, unlike the Gabor approach, it produces features that are strongly aligned with heuristics-based intuition about the data, resulting in qualitatively superior explanations. As such, the features can be tied directly back to the scientific problem and used to adjust the experimental system to produce triangle plots of desirable quality.
It is reasonable to assume that the Gabor technique excels due to its ability to represent the ridged pattern in the triangle region which is considered an important characteristic of a good scan. We see that including a single Gabor filter allows us to recover nearly all of the performance without compromising the model’s interpretability. Our use of a generic global optimizer is also a limiting factor, especially in cases where the synthetic function contains more detailed terms. Alternatives include modifying the objective function to be convex within a feature-informed neighborhood of the global optimum or deriving a custom heuristic optimizer with superior convergence.

### v5

@@ 2.5 Synthetic data modeling approach
In practice, this involves applying an ideal low-pass filter, such as a Gaussian filter. To focus on the details of the structure, we use the magnitude gradient of the low-pass filtered image. While this can be thought of as a simple edge detection, we do not have a priori knowledge about the appropriate scale of the Gaussian for any particular image. The presence of the low-pass filter makes this approach notably more insensitive to certain classes of noise, which is desirable for our use case. To quantify the similarity between an experimentally acquired and synthetic triangle plot, we use the $L^{2}$-norm applied to both transformed images, assuming that the two images are visually similar if their distance in this transformed space is small.
Definition.
@@
$q^{*},\sigma^{*}=\arg\!\min_{q,\sigma}\big{\{}\lambda\,\mathcal{S}_{q,\sigma}^{\nabla_{\Gamma}}(\mathcal{I}_{\rm exp},\mathcal{I}_{\rm sim})+(1-\lambda)\,\mathcal{S}_{q}^{I}(\mathcal{I}_{\rm exp},\mathcal{I}_{\rm sim})+\epsilon(q)\big{\}},$ (6)
where $\mathcal{S}_{\sigma}^{\nabla_{\Gamma}}$ is the similarity measure using the gradient of the Gaussian transform, $\nabla_{\Gamma}(U)=\nabla(\Gamma_{\sigma}(U))$; $\mathcal{S}^{I}$ is the similarity measure using the identity transform, $I(U)=U$; $\lambda$ is a hyperparameter giving us control over the balance between trusting $\mathcal{S}_{\sigma}^{\nabla_{\Gamma}}$ and $\mathcal{S}^{I}$ similarity measures; and
@@
for some arbitrarily large penalty constant $C$. In practice, it is desirable to make this penalty larger than the other terms of the objective function so that the constraints are satisfied.
Since the initial guess for optimization is determined solely based on the raw image, we use the differential evolution global optimizer implemented in SciPy [33] to remedy this limitation. This makes our method less sensitive to initial values than if we had used a local optimizer but incurs a greater computational cost. A local optimizer might still be viable in cases where the phenomenological model is simpler, e.g., with fewer parameters or with guaranteed convexity. In an experimental context, we have access to additional data tied to each image to aid in establishing initial guesses for optimization.
@@ 3 Results
We present the results of experiments with the Gabor filterbank approach, the synthetic data modeling approach, and the hybrid approach combining the synthetic features with the single most informative Gabor feature. The transformation procedures described in sections 2.3 and 2.5 are applied to each experimentally observed triangle to produce the final vector of features $\textbf{v}(\mathcal{I}_{exp})$. In each case, the vectorized features $\textbf{v}(\mathcal{I}_{exp})$ are used to train and test an EBM model. Since the dataset is quite imbalanced in terms of good and bad data, the training is performed using five six-fold stratified cross-validations, with $10~{}\%$ of the data withheld for testing purposes. To determine the optimal $k$-fold configuration, we performed the stratified cross-validation for $k$ ranging between $1$ and $10$, as we found that $k=6$ achieved the most performant set of models.
To improve the accuracy and interpretability at a modest training time cost, we utilize the smoothing and greedy rounds parameters [34]. For a more thorough discussion of this, see Ref. [35]. We carry out five six-fold stratified cross-validations and report averaged results as well as type I and type II errors for each method. For completeness, we also include the averaged confusion matrices.
The results from all experiments, presented in table LABEL:tab:model_comparison, show a relatively close performance for all three methods, with the Gabor fitting approach only slightly outperforming the more intuitive synthetic data modeling approach, at $92.5(1.2)~{}\%$ vs. $90.5(1.0)~{}\%$, respectively. Interestingly, the inclusion of a single, most important Gabor filter in the vector of features obtained using synthetic data modeling brings the EMBs performance back up to $91.7(8)~{}\%$. Importantly, the model resulting from the hybrid vectorization retains the majority of the performance while the features in the hybridized vector remain well aligned with the physical intuition of the problem. This validates that the optimal fit captures the location and average intensity of the walls and diagonal components. Modifying the diagonal component function to capture the ridged pattern observed in the experimental data could further improve the performance.
Figure 3: The feature importance plot for (a) the Gabor filterbank vectorization approach, (b) the synthetic data modeling vectorization approach, and (c) the hybrid vectorization approach. The mean absolute score refers to the extent to which a feature contributes to all predictions. Feature curve for (d) the oriented Gabor feature $\mathcal{G}_{16,16}^{45}$, (e) the fit fitness feature $\mathcal{F}$ of the synthetic triangle fit, and (f) the diagonal boundary component $\mathcal{D}_{B}$ of the synthetic triangle fit for the synthetic data modeling approach shown in orange and for the hybrid approach shown in blue. The $x$-axis indicates the value attained by the feature [normalized to $(0,1)$ in (d) and (e)], and the score indicates the log odds towards good (positive) and bad (negative) classes for each respective feature. The histograms in (g-i) depict the relative class distribution as a function of the normalized feature value, with cyan representing good triangles and pink representing bad triangles.
@@ 3.1 Prediction interpretability
The feature importance plot for the synthetic data modeling approach, shown in figure 3(b), reveals that this model also prioritizes features characterizing the triangle region, with $\mathcal{D}_{B}$ (the diagonal boundary component) and $\mathcal{D}_{M}$ (the diagonal magnitude component) being among the top three filters. The most important feature here is the fit fitness $\mathcal{F}$, quantifying how close the vectorized representation matches the experimental data. Looking at the feature curve for $\mathcal{F}$ depicted in figure 3(e), we see that for a modest cost associated with fitting, that is when $10^{2}\lessapprox\mathcal{F}\lessapprox 10^{3}$, the log odds of belonging to the good class are positive. For bad triangles, the associated cost is either disproportionately low or high. On the low end of the spectrum, this is likely due to the problem of fitting a bad triangle with only two rectangular regions being a significantly simpler optimization problem. On the high end, the exceptionally high cost is generally associated with poor quality of the data.
@@
The second most important feature in the synthetic data modeling approach is $\mathcal{D}_{B}$. The $\mathcal{D}_{B}$ feature curve, shown in orange in figure 4(f), indicates that the log odds that a triangle region belongs to a good plot increase with $\mathcal{D}_{B}$ once it extends past 28. The log odds of a triangle region belonging to a good plot are positive when $8535<\mathcal{D}_{B}<$ 85. The $\mathcal{D}_{B}$ feature quantifies the extent to which the triangle region stretches in the screening gate-screening gate plot, as depicted in figure 2(c). When $\mathcal{D}_{B}<35$, it is likely that the fitted triangle region overlaps with the horizontal and/or vertical sigmoid, and thus $\mathcal{D}_{B}$ would not be a reliable discriminant between good and bad class. While theoretically vector $\mathcal{D}_{B}$ can extend to the bottom left corner in figure 2(c), for large values of $\mathcal{D}_{B}$ the hypotenuse becomes relatively short which makes finding a reliable fit challenging. This dependency is confirmed by the feature plot, indicating that for $\mathcal{D}_{B}>85$, the log odds of a triangle region belonging to a good class fall below 00. The importance of the $\mathcal{D}_{B}$ filter is further evidenced by its dominant role in the hybrid approach, where it occupies the third most important position, see figure 4(c).
@@
It is important to emphasize that, due to the way EBMs are trained, the feature curves are not the same as the marginal distribution of feature values. Rather, each curve is affected by the complete vector of features $\textbf{v}(\mathcal{I}_{exp})$ used to train the model. This means that individual features cannot be treated as entirely independent from one another. This is confirmed by the slight difference between the shape of the $\mathcal{D}_{B}$ curve for the synthetic data modeling approach vs. the hybrid approach, shown in figure 3(f) in orange and blue, respectively.
@@
The hybrid model achieves comparable performance to the Gabor-based model while including only a single Gabor feature in addition to the synthetic features. It thus retains the improved intelligibility of the synthetic features without a significant reduction in accuracy.
@@
Figure 4: (a) A sample experimentally acquired bad scan. (e, i) Two sample experimentally acquired good scans. (b, f, j) The fitted synthetic triangle plots for scans (a), (e), and (i), respectively. (c, g, k) Gabor-filter-based local interpretation of scans is shown in (a), (e), and (i), respectively. (d, h, l) The feature importance for the hybrid vectorization method for scans (a), (e), and (i), respectively. The intercept coefficient $\alpha=4$ is removed from plots (d), (h), and (l) for clarity.
@@
EBMs also give us the ability to examine predictions for individual images. Figure 4(a), (b), and (c) show an example of a bad image, its synthetic fit, and Gabor filter representation, respectively. The synthetic triangle fit effectively captures the structural content of the experimental image, exhibiting no triangle region. The relative feature importance plot, shown in figure 4(d), also agrees with this observation, indicating $\mathcal{V}_{B}$ (the vertical boundary component) and $\mathcal{H}_{B}$ (the horizontal boundary component) as the most dominant features contributing to the negative (i.e., bad class) prediction.
@@
In all three examples, there is a direct observable relationship between the fit parameters and the visual features of the synthetic triangle. In practical applications, these visual features can lead to thresholding decisions and derived quantities for choosing operating points in the voltage space. For example, in the case of figure 4(i), a good operating point might be right inside the triangle region above the diagonal boundary, bisecting the area of the triangle at around $(V_{S_{3}},V_{S_{2}})=(0.09~{}$\mathrm{V}$,0.09~{}$\mathrm{V}$)$ [at point (25,25) in figure 4(j)]. Using the $\mathcal{D}_{B}$ and $\mathcal{D}_{\theta}$ with the knowledge of the boundaries of the image in voltage space, along with the $\mathcal{V}_{B}$ and $\mathcal{H}_{B}$, we can create a normal vector describing the direction in which the triangle is expanding. This normal vector describes the relative strengths of the associated screening gates. Ideally, the normal vector of the triangle would point symmetrically towards the bottom left corner, meaning that both screening gates have equal action on the forming current channel. But, more often than not, this vector tends to vary. By re-scaling the normal vector and defining the magnitude to begin at ($\mathcal{V}_{B}$,$\mathcal{H}_{B}$) with the end pointing to the diagonal boundary of the triangle, the vector can now encode size information. If we then reduce the magnitude by $10~{}\%$, we have an initial guess, inside the bounds of the triangle, for the operating point. We can also use the vector to calculate the area of the triangle in voltage space, which can be thresholded to ensure it is large enough to proceed.
@@ 4 Conclusion and outlook
In this work, we demonstrate an alternative approach to vectorizing image data for use with EBMs that relies on generating synthetic data that best fits the experimentally acquired scans. We show that the new method is better adapted to the complexity of the triangle plots dataset. While the original vectorization method, invoking the Gabor filterbank features, produces a model that performs well with the classification task, its interpretability and usefulness are limited as the resulting model does not comprehend the underlying structure of the data. The alternative method, relying on fitting synthetic triangle plots to the experimental data, retains comparable accuracy, failing in about $2~{}\%$ more cases than the Gabor filterbank approach. However, unlike the Gabor approach, it produces features that are strongly aligned with heuristics-based intuition about the data, resulting in qualitatively superior explanations. As such, the features can be tied directly back to the scientific problem and used to adjust the experimental system to produce triangle plots of desirable quality. Future work might include integrating this analysis directly into an automated real-time QD tuning system.
It is reasonable to assume that the Gabor technique excels due to its ability to represent the ridged pattern in the triangle region which is considered an important characteristic of a good scan. We see that including a single Gabor filter—capturing the presence and prevalence of a diagonal activity—allows us to recover nearly all of the performance without compromising the model’s interpretability. Alternatively, to avoid calculating the Gabor filterbank, this behavior could also be encapsulated within the synthetic approach by representing this pattern with a sinusoidal term in the synthetic triangle model.
The one disadvantage of the synthetic data modeling approach, when weighed against the Gabor filterbank, approach is the greater computational cost of the former. This is primarily due to the reliance on a global optimizer, whereas the Gabor-based approach only requires a small number of fixed filters. To remedy this, alternative similarity functions that would improve convergence and be less reliant on a global optimizer could be explored. However, care must be taken to ensure that the qualitative fitness is not reduced.
Finally, while all data used in this work comes from a quad-QD device with an overlapping gate architecture, we expect these results to be generalizable to an entire subclass of devices in Si/$\text{Si}_{x}\text{Ge}_{1-x}$ with different gate designs [36, 21, 37]. Likewise, other electronic materials that can host gate-defined QDs, such as bi-layer graphene [38] and silicon-MOS structures [39], can benefit from ML-enabled characterization and tuning as proposed in this work. The only requirement is that the tuning procedure must involve confining electrons to a 1D channel with a pair (or pairs) of gates so that other additional gates can then divide that channel into a chain of QDs.
Data availability statement
The data that support the findings of this study are openly available at the following URL/DOI: https://doi. org/10.5281/zenodo.14549897. Figure source files are openly available at the following URL/DOI: https://doi. org/10.5281/zenodo.14589568.
@@ Acknowledgments
Appendix
A.1 Misclassified data: A post-hoc analysis
While the synthetic triangles provide a good representation of the crude features captured in experimental data for the majority of the data, there are cases where the vectorized representation is misclassified by the EBM. There are two types of misclassifications: classifying a triangle plot with a well-pronounced triangle region as bad [false-negative, see figure A1 panels (a)–(c)] and classifying a plot without well-defined triangle region as good [false-positive, see figure A1 panels (d)–(e)].
Figure A1: Sample ambiguous triangle plots. (a)–(c) Good-class triangle plots classified by the model as bad. (d)–(e) Bad-class triangle plots classified by the model as good.
There are several reasons for such misclassifications. In figure A1(a) and (b), the plots are excessively accumulated, which would usually indicate a good sample if the user were to zoom out, and there is not sufficient diagonal activity for the sample to be valid. In figure A1(a) the scan window seems too small for the size of the triangle plot. In figure A1(b), on the other hand, we see the shadow of a second channel influencing the readout. This is a rare breakdown of the model’s assumptions where the channels of the QD device were assumed to be independent. The ohmic potential is typically proportional to the maximum current in the image. With high accumulation, the ohmic potential can drift dynamically depending on multiple channels simultaneously since the measured current takes the path of the lowest resistance. This issue only plagues the upper channel of this device with charge sensors, see figure 1(a), since they are not separated by a gate-like isolation of the quad-QD channel on the bottom of the device. To resolve this with the same model, the experimental setup would need to know exactly when the ohmic potential started drifting and compensate for both halves of the charge sensor to keep exactly half the current flowing out through both channels equally. In figure A1(c), the slope indicating the formation of coulomb blockade is relatively weak but technically visible, which, to an experimentalist, indicates that the triangle plot is good. However, in all three plots, one of the channels seems to be missing, which contributed to them being classified as bad even though the triangle region is present in all of them.
In figure A1(d), there is a substantial region with no current inside the triangle region, which seems to suggest that the channel is not accumulated correctly. In figure A1(e), the device is tuned to the wrong regime for the triangle plot where the blockade is controlled by one screening gate more than the other. This can be seen by the fringes being vertical rather than diagonal. Because plots in figure A1(d) and (e) have erroneous dropout regions, they should be considered bad. However, because the presence of the significant diagonal behavior has been vectorized through the synthetic triangle modeling, the EBM classified these images as good. In the plot depicted in figure A1(f), the strong diagonal and ridged pattern indicates a possibly shocked device. Yet, the diagonal rigid region is incorrectly vectorized as a proper triangle, which again leads to incorrect classification by EBM.
ORCID iDs
[…]